OpenStudy (anonymous):

Identify the definite integral that represents the arc length of the curve y=1/x over the interval [1, 3]

OpenStudy (anonymous):

my calculus is super rusty >.< but isnt the formula for arc length: \[\int\limits_{a}^{b}\sqrt{1+y'}dx\] something like that?

OpenStudy (anonymous):

someone needs to com and verify that for me lol <.<

OpenStudy (anonymous):

i dont like calculus >.<

OpenStudy (saifoo.khan):

Seems Scary! :/

myininaya (myininaya):

that y' should be squared

OpenStudy (anonymous):

i forgot a squared in the sqrt. its \[\int\limits_{a}^{b}\sqrt{1+(y')^2}dx\]

OpenStudy (anonymous):

i noticed >.< thanks! did i mention i dont like calculus yet? >.<

OpenStudy (anonymous):

O it's 1+(1/x^4)?

OpenStudy (anonymous):

thats whats going to go inside the square root :)